### v1

Abstract
Purpose: Surgical phase recognition (SPR) is an integral component of surgical data science, enabling high-level surgical analysis. End-to-end trained neural networks that predict surgical phase directly from videos have shown excellent performance on benchmarks. However, these models struggle with robustness due to non-causal associations in the training set, resulting in poor generalizability. Our goal is to improve model robustness to variations in the surgical videos by leveraging the digital twin (DT) paradigm – an intermediary layer to separate high-level analysis (SPR) from low-level processing (geometric understanding). This approach takes advantage of the recent vision foundation models that ensure reliable low-level scene understanding to craft DT-based scene representations that support various high-level tasks.
Methods: We present a DT-based framework for SPR from videos. The framework employs vision foundation models to extract representations. We embed the representation in place of raw video inputs in the state-of-the-art Surgformer model. The framework is trained on the Cholec80 dataset and evaluated on out-of-distribution (OOD) and corrupted test samples.
Results: Contrary to the vulnerability of the baseline model, our framework demonstrates strong robustness on both OOD and corrupted samples, with a video-level accuracy of 51.1 on the challenging CRCD dataset, 96.0 on an internal robotics training dataset, and 64.4 on a highly corrupted Cholec80 test set.
Conclusion: Our findings lend support to the thesis that DT-based scene representations are effective in enhancing model robustness. Future work will seek to improve the feature informativeness, automate feature extraction, and incorporate interpretability for a more comprehensive framework.
keywords:
surgical video, cholecystectomy, out-of-distribution generalization
1 Introduction
Surgical phase recognition (SPR) is a critical component of surgical data science, providing essential insights for surgical workflow analysis, skill assessment, and numerous other applications. With the rise of deep learning, end-to-end feed-forward networks [1, 2, 3, 4, 5, 6] have demonstrated strong performance on SPR benchmarks [1, 7]. However, concerns around the robustness of deep learning models [8] are growing, and are of particular importance in high-stakes domains, such as surgical data science. Robustness is vital for clinical translation [9], but end-to-end trained feed-forward networks – that in recent history have performed best in many complex tasks – often exhibit low robustness due to various non-causal associations in the training set, which is confounded by subtle factors in high-dimensional data such as images and videos.
The digital twin (DT) paradigm, as illustrated in Figure 1, constructs and maintains computational representations of real-world environments. It serves as an intermediary layer between the video pixels and SPR, achieving a separation between high-level analysis and low-level processing. From the real-world observations (physical twin) with high variance, DT applies semantic identification and geometric reconstruction techniques such as segmentation, depth estimation, 3D reconstruction, and pose estimation to extract a digital counterpart, a DT-based scene representation. High-level analysis is then trained and conducted on this low variance representation to reduce non-causal learning when training feed-forward networks, thereby enhancing the model’s robustness. Nowadays, DT receives increasing attention, and extracting DT-based scene representations becomes more feasible due to the emergence of vision foundation models for low-level processing, like the segment anything model 2 (SAM2)[10] and DepthAnything[11]. These models have shown impressive zero-shot generalization and robustness [12], thanks to their large model capacity, vast training datasets, and advanced self- or semi-supervised learning mechanisms.
As an instantiation, we propose a robust SPR framework via DT-based scene representation. We harness the strong performance and robustness of vision foundation models by utilizing SAM2 [10] and DepthAnything [11] to extract basic DT-based scene representations. These structured representations then replace video inputs in Surgformer [6] model enabling SPR from surgical videos via the DT with enhanced robustness. We train our framework based on the Cholec80 [1] dataset and evaluate its performance and robustness on both original and corrupted images from the Cholec80 test dataset as well as the CRCD [13] and an internal cholecystectomy dataset acquired using the da Vinci Research Kit (dVRK) on ex vivo porcine specimens as OOD test samples. Our results show that our framework demonstrates superior robustness against image corruptions and domain shifts where the end-to-end trained baseline models fail. These findings support the effectiveness of using DT-based scene representations to enhance model robustness, paving the way for more reliable high-level analysis for surgical data and accelerating the clinical translation of research outcomes.
Our contribution can be summarized as follows: (1) A robust SPR framework from surgical video via a DT-based scene representation; (2) An approach to form and apply a DT-based scene representation to state-of-the-art models; and (3) Validation of the robustness of the proposed framework on corrupted and OOD test data.
Figure 1: Illustration of the digital twin paradigm.
2 Related Work
2.1 Surgical Phase Recognition
Early works [14, 15, 16, 17] for SPR mainly applied statistical models such as hidden Markov models (HMMs) [17, 16, 15]. They formulated the procedure as a finite-state machine and represented the HMMs as transition relations between phases and observations. Due to the weak hand-crafted visual signals [16], to achieve more satisfying performance, some methods [15, 14, 16] relied on additional sensors. With the advancement of deep learning, neural network-based methods became the state-of-the-art method for surgical phase recognition. EndoNet [1] started to apply convolutional neural networks (CNN) for feature extraction and use the HMMs for temporal integration. Some methods [2] used recurrent models to leverage spatial and temporal information and achieved great improvement. Applying temporal convolutional networks (TCNs) [18] was another design choice due to the vanishing gradient issue of the recurrent models. The transformer [19] network and ViT [20] introduce the attention mechanism to deal with the longer sequence with parallel training. Trans-SVNet [4] applied the attention mechanism to process the spatial and temporal features. Killeen et al.  [21] proposed the Pelphix architecture to predict the surgical phase from X-ray images. SKiT [5] encode image features with ViT [20] and hierarchically decode features using transformer-based basic layers. Surgformer [6] proposed divided spatial-temporal attention and takes a limited set of sparse frames as input to improve model performance.
2.2 Robustness in Surgical Data Science
Despite the excellent performance, the end-to-end learning fashion for those feed-forward networks (FFNs) poses the challenge of model robustness [8] when facing domain gap or non-adversarial corruptions [22]. The lack of reliability of the proposed algorithms hindered the clinical translation for the research achievement for surgical data science [9, 23]. Researchers started to take low-level vision tasks like instrument segmentation and assess the models’ robustness by adding corruptions to the input image [24]. Ding et al. added non-adversarial physical corruptions to the test data [22, 25, 26] taking advantage of the precise replayability of the da Vinci Research Kit [27]. However, to our best knowledge, there is a lack of assessment of the model’s robustness for high-level analysis of surgical procedures like surgical phase recognition.
2.3 Digital Twin-based Scene Representation
The DT paradigm separates the low-level processing from high-level analysis [23]. The DT-based scene representation generated from low-level processing forms a universal representation to filter the non-causal features for the high-level analysis. Previous works [28, 29, 30, 31, 32] focus on the extraction of the DT-based scene representation from external tracking devices and markers or visual input. Ding et al. [32] applied the DT-based scene representation in an embodied surgical system and improved the robustness of the surgical automation, providing evidence of the effectiveness of the DT-based scene representation in improving the robustness of the downstream application.
3 Methods
Figure 2 presents an overview of our proposed robust surgical phase recognition framework using digital twin (DT)-based scene representations. The framework comprises three primary modules: representation extraction, DT-based patch embedding, and DT-based SPR. In the representation extraction module, raw visual inputs, such as surgical videos, are processed using vision foundation models to extract fundamental representations like depth maps and segmentation masks. These representations are further processed in the DT-based patch embedding module to form a latent representation. The final module, DT-based SPR, applies the state-of-the-art SPR model (Surgformer), using the embedded representations for training and inference. Each of these modules will be detailed in the following subsections.
Figure 2: Illustration of the surgical phase recognition framework via DT-based scene representation.
3.1 Representation Extraction
For representation extraction, we employ SAM2 [10] for instance segmentation and DepthAnything [11] for monocular depth estimation.
3.1.1 Instance segmentation
We define 10 classes, as outlined in Table 1. These include one target tissue (gallbladder), one specimen bag, and six types of surgical instruments: grasper, bipolar, hook, scissors, clipper, and irrigator. To distinguish between multiple graspers that appear simultaneously, we differentiate them based on their insertion positions.
Table 1: Instance list for instance segmentation. Each pair is the instance name and id.
gallbladder 1 left grasper 2 top grasper 3 right grasper 4 bipolar 5
hook 6 scissors 7 clipper 8 irrigator 9 specimen bag 10
[…]
00footnotetext:
For instance segmentation, we use SAM2 [10], providing a point prompt for each instance at the frame where it first appears. SAM2 utilizes a Vision Transformer (ViT) [20] as the image encoder and incorporates a memory attention module to integrate spatial and temporal information from previous and prompted frames. The point prompt is encoded using a prompt encoder, and a mask decoder processes the encoded prompt and image features to generate mask predictions while updating memory.
3.1.2 Depth estimation
We utilize the DepthAnything [11] model to predict depth directly from a single frame. The model follows an encoder-decoder architecture, estimating relative disparity from monocular images. The predicted disparity is normalized to a 0-1 scale, which we use to represent relative depth.
3.2 DT-based Patch Embedding
For each frame, the raw representation consists of 10 binary segmentation masks for each class and a normalized disparity map. We format the segmentation masks into a 10-channel, one-hot encoded tensor, then concatenate the disparity map as the 11th channel. We apply a patch embedding mechanism across a sequence of frames, transforming the tensor into a spatial-temporal sequence of embeddings. This sequence is formatted to suit the input requirements of transformer-based encoders [19].
3.3 DT-based SPR
We use Surgformer [6] as the prediction model to determine the surgical phase for each frame. Surgformer processes a video clip of fixed length and sample rate, adding a classification embedding to aggregate features for phase prediction. It integrates hierarchical temporal attention (HTA) into the TimesFormer [33] encoder to capture both global and local temporal information. The processed classification embedding from the final layer is passed to a linear prediction head, predicting the phase of the last frame in online prediction settings, we replace the raw video clip’s patch embedding with our DT-based patch embedding to form the DT-based SPR.
4 Experiment
In this section, we describe the experiments conducted to assess the performance and robustness of our proposed method. We present three key experiments: out-of-distribution (OOD) generalization, robustness against image corruptions, and ablation studies. For all experiments, both our method and the baseline model are trained on the public Cholec80 dataset [1]. In the OOD generalization experiment, we evaluate our method on the original Cholec80 test set, as well as the CRCD [13] and an internal robotics training dataset, serving as OOD test sets. To assess robustness against image corruption, we apply both combined and individual image corruption to examine the model’s robustness. The ablation study explores the contributions and effectiveness of the low-level representation and design choices of DT-based patch embedding. The following subsections will outline the experimental setup, including datasets, evaluation metrics, and model configurations, followed by the results of these three experiments.
4.1 Experiment Setting
Figure 3: Visual example of the original and corrupted Cholec80 [1], CRCD [13],and the robotics training dataset.
4.1.1 Dataset
We train our method and baselines on the Cholec80 dataset [1], which contains 80 cholecystectomy videos, each annotated with 7 surgical phases. The dataset is evenly split into training/validation and testing sets. For efficiency, we downsample the videos to 1 frame per second (fps). Due to the requirement of point prompts, we annotate and use videos 1-10 from the original training split for training and videos 41-50 from the original test split for testing. We also apply digital corruptions like Gaussian noise to these test samples to form a corrupted test set to evaluate models’ robustness to digital corruptions. The CRCD [13] dataset includes The internal robotics training dataset consists of 10 videos capturing robotic ex-vivo cholecystectomy training on the dVRK platform, specifically performing cutting and clipping tasks.
4.1.2 Evaluation Metrics
In this work, we evaluate performance using four key metrics: video-level accuracy, phase-level precision, recall, and Jaccard index. Video-level accuracy represents the mean percentage of correctly recognized frames for each video, with both the mean and standard deviation reported across videos. Given the class imbalance, phase-level metrics offer further insights. Precision is defined as the ratio of true positives to total predictions, recall is the ratio of true positives to actual instances, and the Jaccard index measures the overlap between predictions and ground truths. For each phase, we collect the total number of true positives, false positives, and false negatives, and compute precision, recall, and Jaccard. Finally, we calculate the mean for these metrics across all phases. For the CRCD [13] and robotics training test data, since there only exists one phase for each dataset, we only report recall for phase-level metrics.
4.1.3 Model Setting
We apply the ViT-Large encoder for both the SAM2 [10] model and depthAnything [11] model and use the officially provided pre-trained model for inference. For the Surgformer [6] model, we apply the official implementation for online SPR. We follow their settings and apply the input frame number of 16, the sampling rate of 4, and the learning rate of 0.0005. We train each model for 20 epochs with a batch size of 4.
4.2 OOD Generalization
As shown in Figure 3, although these datasets depict the same procedure, there is a notable appearance gap between the Cholec80 [1], CRCD [13], and robotic cholecystectomy datasets, making them well-suited for OOD evaluation. The results, presented in the OOD section of Table 2, highlight that while the Surgformer baseline performs strongly on the Cholec80 test data, it struggles significantly on both OOD test sets. In contrast, our model demonstrates reasonable results on the CRCD [13] and robotics training datasets, underscoring its robustness to OOD variations.
4.3 Robustness against Corruption
We apply hue transformation, brightness adjustment, and contrast alteration as individual image corruptions, and combine them to create the combined corruption. Examples of these corrupted images are shown in Figure 3. As seen in the corruption results in Table 2, the baseline models experience a significant performance drop under these corruptions. In contrast, our framework demonstrates strong robustness overall, with only a slight decline in performance, which is primarily due to a less accurate segmentation mask and depth predictions caused by the corruptions.
Table 2: Experiment results
OOD Cholec80 [1] - Clean CRCD [13] Robitics Training
Methods Acc Prec Rec Jac Acc Rec Acc Rec
[…]
Corruption Cholec80 [1] - Combination Cholec80 [1] - Hue
Methods Acc Prec Rec Jac Acc Prec Rec Jac
[…]
Corruption Cholec80 [1] - Brightness Cholec80 [1] - Contrast
Methods Acc Prec Rec Jac Acc Prec Rec Jac
[…]
00footnotetext: Ours-d, ours-s, ours-c, and ours mean depth-only, segmentation-only, colorized-mask, and final model.
4.4 Ablation Study
The ablation study examines the impact of incorporating depth and segmentation information, as well as our choices for embedding representation. To assess the contribution of depth and segmentation, we create depth-only and segmentation-only models by removing one modality at a time from the input. The results, presented in Table 2, highlight the performance differences. Additionally, we compare our embedding design with an alternative approach that uses colorized masks instead of one-hot encoding.
Compared to the baseline model, all variants demonstrate improved performance on out-of-distribution (OOD) test samples, validating the efficacy of enhancing robustness through a DT-based scene representation. Among the variants, the segmentation-only model exhibits relatively strong robustness but performs slightly worse across all test samples without the benefit of depth guidance. In contrast, the depth-only model achieves better benchmark performance but shows reduced robustness, particularly on heavily corrupted images and OOD samples, compared to the final model. These findings suggest that depth provides valuable complementary details to instance segmentation but, when used in isolation, may introduce vulnerability. The depthAnything [11] model also exhibits limitations in handling image corruption.
The colorized mask embedding approach behaves similarly to the depth-only model, offering stronger overall performance but reduced robustness compared to the final model. In contrast, the final model, which more closely mirrors the behavior of the segmentation-only variant, demonstrates that one-hot encoded embeddings ensure segmentation contributes more to the final prediction, ensuring the model’s robustness.
5 Discussion
5.1 Inspiration
Our DT-based SPR framework exhibits strong robustness in high-level surgical analysis tasks, supporting the efficacy of the DT paradigm as an intermediary layer for enhancing surgical data science research. The advent of vision foundation models enables a highly generalizable and reliable low-level processing pipeline, facilitating the extraction and formation of DT-based scene representations. We anticipate that DT-based research will attract attention for its potential to develop more generalizable and interpretable systems in surgical data science, potentially accelerating the clinical translation of these research advancements.
5.2 Limitation & Future Work
As a proof of concept and an initial exploration of DT-based high-level analysis, further development is required to create a more comprehensive and practical framework for SPR and other high-level analytical tasks. First, in our current framework, we leverage SAM2, which to distinguish semantic instrument classes, requires minimal interaction. Overcoming this limitation is easily possible through recent advances in open-vocabulary vision foundation models [34] or vision-language models that can automatically generate this interaction, presenting a promising direction for future work. Second, despite using the same model capacity, on clean data, and under some corruptions our framework’s benchmark performance lags behind state-of-the-art methods, indicating that it may not yet extract sufficiently rich information for SPR. Further exploration into additional low-level features, such as scene flow, which can provide informative data for SPR and be reliably obtained, will be essential for improving performance. Third, while this work focuses on model robustness, we have not yet explored the potential interpretability offered by the DT-based scene representation which is relevant to clinical translation, as it builds trust by clarifying the decision-making process, an avenue we seek to further explore in future efforts.
6 Conclusion
We propose a surgical video analysis framework using a DT-based scene representation. This framework demonstrates strong robustness against OOD and corrupted test samples, validating the effectiveness of high-level surgical analysis via DT-based scene representation and the reliability of low-level representation extraction through vision foundation models. Our work highlights a promising research direction for developing more robust algorithms in surgical data science, moving beyond recent end-to-end deep learning approaches. To create a more comprehensive framework, future research should focus on enhancing the informativeness of DT-based feature representations and improving the efficacy of the feature extraction pipeline. Additionally, incorporating explainable AI techniques to enhance interpretability is likely beneficial for facilitating clinical translation.
[…]
Acknowledgements This research is supported in part by a collaborative research agreement with the MultiScale Medical Robotics Center at The Chinese University of Hong Kong and a Seed Grant from the Malone Center for Engineering in Healthcare.